3,400 principle earning 3% compound after 5 years find the balance in the account...?
i need help with this problem.......can u please help me.i'm an high school student that needs your help!

Answer by stanbon(75887)

3,400 principle earning 3% compound after 5 years find the balance in the account...?
---------------
A(t) = P(1+(r/n))^(nt)
A(5) = 3400(1+(0.03)/1)^(1*5)
A(5) = 3400(1.03)^5
A(5) = 3400*1.159274074..
A(5) = $3941.53
==================
Cheers,
Stan H.